An interesting (but oversimplified) model of an at
Answer #1

In the give situation the force of attraction due to proton will cause the centripital acceleration so the following equation can be written:

\large \frac{mv^2}{r} = \frac{ke^2}{r^2}

So after solving we get,

\large v = \sqrt{\left(\frac{ke^2}{mr} \right )} = 1395561.31 = 1.396*10^6 m/s